Exercise 15-08

Buffalo Inc. recently hired a new accountant with extensive experience in accounting for partnerships. Because of the pressure of the new job, the accountant was unable to review what he had learned earlier about corporation accounting. During the first month, he made the following entries for the corporation’s capital stock.

Date

Account Titles and Explanation

Debit

Credit

May 2

Cash

204,000

   Capital Stock

204,000

(Issued 12,000 shares of $5 par value common stock at $17 per share)

May 10

Cash

800,000

   Capital Stock

800,000

(Issued 10,000 shares of $40 par value preferred stock at $80 per share)

May 15

Capital Stock

14,400

   Cash

14,400

(Purchased 900 shares of common stock for the treasury at $16 per share)

May 31

Cash

11,200

   Capital Stock

5,600

   Gain on Sale of Stock

5,600

(Sold 560 shares of treasury stock at $20 per share)


On the basis of the explanation for each entry, prepare the entries that should have been made for the capital stock transactions. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

May 2May 10May 15May 31

May 2May 10May 15May 31

May 2May 10May 15May 31

May 2May 10May 15May 31

SHOW LIST OF ACCOUNTS

LINK TO TEXT

LINK TO TEXT

Question Attempts: 0 of 3 used

SAVE FOR LATER

SUBMIT ANSWER
Answer #1
Date Account Titles and Explanation Debit Credit
May 2 Cash 204000
       Common Stock 60000 =12000*5
       Paid in Capital in excess of par-Common Stock 144000
May 10 Cash 800000
       Preferred Stock 400000 =10000*40
       Paid in Capital in excess of par-Preferred Stock 400000
May 15 Treasury Stock 14400
       Cash 14400
May 31 Cash 11200
       Treasury Stock 8960 =560*16
       Paid-in Capital from Treasury Stock 2240